OpenStudy (anonymous):

What is the Gcf of 28 and 84

OpenStudy (anonymous):

Do you know what GCF mean?

OpenStudy (anonymous):

28 would be the answer.

OpenStudy (dangerousjesse):

Direct answer.. Ouch

OpenStudy (dangerousjesse):

Find the divisors of each integer and select the largest element they have in common: The divisors of 28 are: \(1, 2, 4, 7, 14, 28\) The divisors of 84 are: \(1, 2, 3, 4, 6, 7, 12, 14, 21, 28, 42, 84\) The largest number common to both divisor lists is 28: divisors of \(28: 1, 2, 4, 7, 14, 28\) divisors of \(84: 1, 2, 3, 4, 6, 7, 12, 14, 21, 28, 42, 84\)

OpenStudy (dangerousjesse):

So in this case, GCD=GCF :)

OpenStudy (anonymous):

28 = 2 * 2*7 84 = 2*2*3*7 Since 7 is the common factor of both the numbers hence 7 is the GCD of 28 and 84

OpenStudy (ikram002p):

you can use several methods :- 1- find all common factors of 28 and 84 and find the greatest 2_ use Eclids algorithm 3_ GCD(a,b)=axb/ LCF(a,b)